# Characterizations of amenability for noncommutative dynamical systems and Fell bundles

### v1

@@ 5.1. Reductions to open subgroups and full cross-sectional $\mathrm{C}^{*}$-algebras
Assume $H$ is open in $G$ and think $C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B})$ by extending each $f\in C_{c}(\mathcal{B}_{H})$ as 0 outside $H.$ This gives $L^{1}(\mathcal{B}_{H})\subseteq L^{1}(\mathcal{B})$ and, passing to enveloping $\mathrm{C}^{*}$-algebras, we get a *-homomorphism $\iota\colon C^{*}(\mathcal{B}_{H})\to C^{*}(\mathcal{B})$. To prove $\iota$ is injective we take a nondegenerate *-representation $T\colon C^{*}(\mathcal{B}_{H})\to\mathbb{B}(X)$ with faithful integrated form $\tilde{T}\colon C^{*}(\mathcal{B}_{H})\to\mathbb{B}(X).$ By [17, Theorem 2.1], $T$ is $\mathcal{B}$-positive and it can be induced to a *-representation $S\colon\mathcal{B}\to\mathbb{B}(Y)$ by the process of [16, Ch. XI]. Since $G/H$ is discrete, we can use [16, XI 14.21] to deduce that $T$ is a subrepresentation of $S|_{\mathcal{B}_{H}}$. Consequently, $\tilde{T}|_{C_{c}(\mathcal{B}_{H})}$ is a subrepresentation of $\tilde{S}\circ\iota|_{C_{c}(\mathcal{B}_{H})}$ and this implies that $\iota$ is an isometry because, for all $f\in C_{c}(\mathcal{B}_{H}),$ $\|\iota(f)\|\leq\|f\|=\|\tilde{T}_{f}\|\leq\|\tilde{S}_{\iota(f)}\|\leq\|\iota(f)\|.$ From now on we think $C^{*}(\mathcal{B}_{H})\subseteq C^{*}(\mathcal{B})$ and omit any mention of $\iota$.
Notice that for all $f,g\in C_{c}(\mathcal{B}_{H})$ and $h\in C_{c}(\mathcal{B})$ we have $f*h*g=f*h|_{H}*g.$ Since $C_{c}(\mathcal{B}_{H})$ contains an approximate unit of $C^{*}(\mathcal{B}_{H}),$ the restriction map $E_{c}\colon C_{c}(\mathcal{B})\to C_{c}(\mathcal{B}_{H}),$ $E_{c}(h):=h|_{H},$ is a contraction with respect to the $\mathrm{C}^{*}$-norm. By Tomiyama’s Theorem [10, Theorem 1.5.10], the continuous extension $E\colon C^{*}(\mathcal{B})\to C^{*}(\mathcal{B}_{H})$ of $E_{c}$ is a conditional expectation.
@@ 5.2. Reductions to open subgroups and reduced cross-sectional $\mathrm{C}^{*}$-algebras
We continue assuming that $H$ is open in $G$ and denote $\lambda^{H}$ and $\lambda^{G}$ the respective left regular representations. To prove that $C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B})$ induces and inclusion of $\mathrm{C}^{*}$-algebras $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})\subseteq C^{*}_{\mathop{\rm r}}(\mathcal{B})$ we take a nondegenerate *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$ with $T|_{B_{e}}$ faithful. By [15, Corollary 2.15], there exists a faithful *-representation $\pi\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to\mathbb{B}(L^{2}(G,X))$ such that, for all $f\in C_{c}(\mathcal{B}),$ $\pi(f)=[\lambda^{G}\tilde{\otimes}T]_{f}.$ For all $f\in C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B}),$ $\pi(f)=[\lambda^{G}|_{H}\tilde{\otimes}T|_{\mathcal{B}_{H}}]_{f}$ and $\lambda^{G}|_{H}$ is a direct sum of copies of $\lambda^{H}$ because $G$ is the disjoint union of open cosets $Ht$. Then $\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B})}=\|[\lambda^{G}\tilde{\otimes}T]_{f}\|=\|[\lambda^{H}\tilde{\otimes}T|_{\mathcal{B}_{H}}]_{f}\|=\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})},$ which implies that we may identify $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ with the closure of $C_{c}(\mathcal{B}_{H})$ in $C^{*}_{\mathop{\rm r}}(\mathcal{B})$.
@@ 5.3. The universal completion of the partial cross-sectional bundle
Assume that $H$ is a closed and normal subgroup of $G$. Let $\mathcal{L}=\{L^{1}(\mathcal{B}_{u})\}_{u\in G/H}$ be the partial cross-sectional bundle over $G/H$ derived from $\mathcal{B}$, and ${\mathcal{B}}_{/H}=\{\dot{B}_{u}\}_{u\in G/H}$ the bundle $\mathrm{C}^{*}$-completion of $\mathcal{L}$, see [16, VIII 6 & VIII 16.7]. By construction, each fiber $L^{1}(\mathcal{B}_{u})$ is the $\|\ \|_{1}$-completion of $C_{c}(\mathcal{B}_{u})$, where $\mathcal{B}_{u}=\{B_{t}\}_{t\in u}$ is the reduction of $\mathcal{B}$ to the coset $u\in G/H$. The natural map $C_{c}(\mathcal{B})\to C_{c}(\mathcal{L}),$ $f\mapsto\tilde{f}$ with $\tilde{f}(u)=f|_{u}$, extends to a *-isomorphim $\Phi\colon L^{1}(\mathcal{B})\to L^{1}(\mathcal{L})$ (see [16, VIII 6.7]). The cross-sectional $\mathrm{C}^{*}$-algebras of $\mathcal{B}$ and $\mathcal{L}$ are the envelping $\mathrm{C}^{*}$-algebras of $L^{1}(\mathcal{B})$ and $L^{1}(\mathcal{L})$, so we may think $C^{*}(\mathcal{B})=C^{*}(\mathcal{L})$. In addition, the map $\rho\colon\mathcal{L}\to{\mathcal{B}}_{/H}$ of [16, VIII 17.7] is a dense embedding in the sense of [16, XI 12.6], so $\{\rho\circ f\colon f\in C_{c}(\mathcal{L})\}$ is dense in $C_{c}({\mathcal{B}}_{/H})$ with respect to the inductive limit topology (use [16, II 14.6]). The construction of the bundle $\mathrm{C}^{*}$-completion implies that all *-representations of $\mathcal{L}$ have the form $T\circ\rho$ for a *-representation $T$ of ${\mathcal{B}}_{/H}.$ Then the continuous *-homomorphism $L^{1}(\mathcal{L})\to L^{1}({\mathcal{B}}_{/H})$ mapping $f\in C_{c}(\mathcal{L})$ to $\rho\circ f$ induces an isomorphism between $C^{*}(\mathcal{L})$ and $C^{*}({\mathcal{B}}_{/H})$. This yields $C^{*}(\mathcal{B})=C^{*}({\mathcal{B}}_{/H})$ or, more precisely, there exists an isomorphism
@@
We claim that the fiber $\dot{B}_{H}$ of ${\mathcal{B}}_{/H}$ over the unit $H$ of $G/H$ is $C^{*}(\mathcal{B}_{H})$. For the bundle $\mathcal{L}$, [16, XI 11.6] says that $C^{*}(\mathcal{B}_{H})=\dot{B}_{H}$ if and only if every *-representation of $L^{1}(\mathcal{B}_{H})$ is $\mathcal{L}$-positive. Using [16, XI 12.7] we get that this last claim is equivalent to say that every *-representation of $\mathcal{B}_{H}$ is $\mathcal{B}$-positive, which is the case [17, Theorem 2.1].
@@ 5.4. The reduced completion of the partial cross-sectional bundle
We want to construct a reduced version ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ of ${\mathcal{B}}_{/H}$ so that $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is isomophic to $C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$ ($H$ is normal and closed in $G$). To do this we take a nondegenerate *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$ with $\phi:=T|_{B_{e}}$ faithful, and let $S\colon\mathcal{L}\to\mathbb{B}(Y)$ be the *-representation corresponding to $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)$ via the process of [16, VIII 15.9]. Using $S$ and the construction of [16, VIII 16.7] we can produce a Fell bundle ${\mathcal{B}}^{\mathop{\rm r}}_{/H}=\{\dot{B}^{\mathop{\rm r}}_{u}\}_{u\in G/H}$ and a dense embedding $\rho^{\mathop{\rm r}}\colon\mathcal{L}\to{\mathcal{B}}^{\mathop{\rm r}}_{/H}$ such that, for all $f\in\mathcal{L},$ $\|\rho^{\mathop{\rm r}}(f)\|=\|S_{f}\|$. By construction, for all $r\in G$; $f\in C_{c}(\mathcal{B}_{rH})$ and $\xi\in Y,$ $S_{f}\xi=\int_{N}\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)_{f(rt)}\xi dt$. Notice that we may replace $f(rt)$ with $\tilde{f}(rN)(rt)$. This ultimately implies that $\tilde{S}\circ\Phi|_{C_{c}(\mathcal{B})}=\widetilde{\operatorname{Ind}}_{B_{e}\uparrow\mathcal{B}}(\phi)|_{C_{c}(\mathcal{B})}$.
The fiber $\dot{B}^{\mathop{\rm r}}_{H}$ of ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ over the unit $H$ of $G/H$ is the completion of $L^{1}(\mathcal{B}_{H})$ with respect to the seminorm $f\mapsto\|S_{f}\|.$ Hence, to prove $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ it suffices to show that for all $f\in C_{c}(\mathcal{B}_{H})$ we have $\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})}=\|S_{f}\|.$ By Fell’s absorption principle [15, 17], $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)\equiv\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(T|_{B_{e}})$ is weakly equivalent ($\sim$) to $\lambda^{G}\otimes T.$ The process of restricting *-representations is continuous with respect to the regional topology [16, VIII 21.20] and, by definition, it is a homeomorphism with respect to the operation of taking integrated forms. This implies
@@
where the last equivalence holds because $\lambda^{G}|_{H}\sim\lambda^{H}$ and because the process of taking tensor products with representations of the base group is continuous with respect to the regional topology. By [15, Corollary 2.15], for all $f\in L^{1}(\mathcal{B}_{H})$ we have $\|S_{f}\|=\|\lambda^{H}\tilde{\otimes}(T|_{\mathcal{B}_{H}})(f)\|=\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})}.$ Then $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$.
@@
The integrated form of $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}_{H}}(T|_{B_{e}})\sim\lambda^{H}\otimes(T|_{\mathcal{B}_{H}})$ is a *-representation $R$ of $L^{1}(\mathcal{B}_{H})$ for which there exists a faithful *-representation $R^{\prime}$ of $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ such that $R^{\prime}\circ\rho|_{L^{1}(\mathcal{B}_{H})}=R,$ see [15, Corollary 2.15]. By [16, XI 12.6-12.7], $S=\operatorname{Ind}_{L^{1}(\mathcal{B}_{H})\uparrow\mathcal{L}}(R)$ and $\operatorname{Ind}_{\dot{B}^{\mathop{\rm r}}_{H}\uparrow{\mathcal{B}}^{\mathop{\rm r}}_{/H}}(R^{\prime})\circ\rho^{\mathop{\rm r}}=\operatorname{Ind}_{L^{1}(\mathcal{B}_{H})\uparrow\mathcal{L}}(R).$ From Proposition 3.1 of [17] we get the existence of a faithful *-representation
@@
The function $\Omega^{\mathop{\rm r}}\colon C_{c}(\mathcal{B})\to C_{c}({\mathcal{B}}^{\mathop{\rm r}}_{/H}),$ $f\mapsto\rho^{\mathop{\rm r}}\circ\tilde{f}$, is a *-homomorphism because $\nu\circ\Omega^{\mathop{\rm r}}=\mu$. Using [16, II 14.6] one can show that the image of $\Omega^{\mathop{\rm r}}$ is dense with respect to the inductive limit topology, so $\Omega^{\mathop{\rm r}}(C_{c}(\mathcal{B}))$ is dense in $C^{*}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$ and $\Omega^{\mathop{\rm r}}$ extends to an isomorphism

### v2

@@ 5.1. Reductions to open subgroups and full cross-sectional $\mathrm{C}^{*}$-algebras
Assume $H$ is open in $G$ and think $C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B})$ by extending each $f\in C_{c}(\mathcal{B}_{H})$ as 0 outside $H.$ This gives $L^{1}(\mathcal{B}_{H})\subseteq L^{1}(\mathcal{B})$ and, passing to enveloping $\mathrm{C}^{*}$-algebras, we get a *-homomorphism $\iota\colon C^{*}(\mathcal{B}_{H})\to C^{*}(\mathcal{B})$. To prove $\iota$ is injective we take a nondegenerate *-representation $T\colon C^{*}(\mathcal{B}_{H})\to\mathbb{B}(X)$ with faithful integrated form $\tilde{T}\colon C^{*}(\mathcal{B}_{H})\to\mathbb{B}(X).$ By [18, Theorem 2.1], $T$ is $\mathcal{B}$-positive and it can be induced to a *-representation $S\colon\mathcal{B}\to\mathbb{B}(Y)$ by the process of [17, Ch. XI]. Since $G/H$ is discrete, we can use [17, XI 14.21] to deduce that $T$ is a subrepresentation of $S|_{\mathcal{B}_{H}}$. Consequently, $\tilde{T}|_{C_{c}(\mathcal{B}_{H})}$ is a subrepresentation of $\tilde{S}\circ\iota|_{C_{c}(\mathcal{B}_{H})}$ and this implies that $\iota$ is an isometry because, for all $f\in C_{c}(\mathcal{B}_{H}),$ $\|\iota(f)\|\leq\|f\|=\|\tilde{T}_{f}\|\leq\|\tilde{S}_{\iota(f)}\|\leq\|\iota(f)\|.$ From now on we think $C^{*}(\mathcal{B}_{H})\subseteq C^{*}(\mathcal{B})$ and omit any mention of $\iota$.
Notice that for all $f,g\in C_{c}(\mathcal{B}_{H})$ and $h\in C_{c}(\mathcal{B})$ we have $f*h*g=f*h|_{H}*g.$ Since $C_{c}(\mathcal{B}_{H})$ contains an approximate unit of $C^{*}(\mathcal{B}_{H}),$ the restriction map $E_{c}\colon C_{c}(\mathcal{B})\to C_{c}(\mathcal{B}_{H}),$ $E_{c}(h):=h|_{H},$ is a contraction with respect to the $\mathrm{C}^{*}$-norm. By Tomiyama’s Theorem [11, Theorem 1.5.10], the continuous extension $E\colon C^{*}(\mathcal{B})\to C^{*}(\mathcal{B}_{H})$ of $E_{c}$ is a conditional expectation.
@@ 5.2. Reductions to open subgroups and reduced cross-sectional $\mathrm{C}^{*}$-algebras
We continue assuming that $H$ is open in $G$ and denote $\lambda^{H}$ and $\lambda^{G}$ the respective left regular representations. To prove that $C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B})$ induces and inclusion of $\mathrm{C}^{*}$-algebras $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})\subseteq C^{*}_{\mathop{\rm r}}(\mathcal{B})$ we take a nondegenerate *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$ with $T|_{B_{e}}$ faithful. By [16, Corollary 2.15], there exists a faithful *-representation $\pi\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to\mathbb{B}(L^{2}(G,X))$ such that, for all $f\in C_{c}(\mathcal{B}),$ $\pi(f)=[\lambda^{G}\tilde{\otimes}T]_{f}.$ For all $f\in C_{c}(\mathcal{B}_{H})\subseteq C_{c}(\mathcal{B}),$ $\pi(f)=[\lambda^{G}|_{H}\tilde{\otimes}T|_{\mathcal{B}_{H}}]_{f}$ and $\lambda^{G}|_{H}$ is a direct sum of copies of $\lambda^{H}$ because $G$ is the disjoint union of open cosets $Ht$. Then $\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B})}=\|[\lambda^{G}\tilde{\otimes}T]_{f}\|=\|[\lambda^{H}\tilde{\otimes}T|_{\mathcal{B}_{H}}]_{f}\|=\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})},$ which implies that we may identify $C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ with the closure of $C_{c}(\mathcal{B}_{H})$ in $C^{*}_{\mathop{\rm r}}(\mathcal{B})$.
@@ 5.3. The universal completion of the partial cross-sectional bundle
Assume that $H$ is a closed and normal subgroup of $G$. Let $\mathcal{L}=\{L^{1}(\mathcal{B}_{u})\}_{u\in G/H}$ be the partial cross-sectional bundle over $G/H$ derived from $\mathcal{B}$, and ${\mathcal{B}}_{/H}=\{\dot{B}_{u}\}_{u\in G/H}$ the bundle $\mathrm{C}^{*}$-completion of $\mathcal{L}$, see [17, VIII 6 & VIII 16.7]. By construction, each fiber $L^{1}(\mathcal{B}_{u})$ is the $\|\ \|_{1}$-completion of $C_{c}(\mathcal{B}_{u})$, where $\mathcal{B}_{u}=\{B_{t}\}_{t\in u}$ is the reduction of $\mathcal{B}$ to the coset $u\in G/H$. The natural map $C_{c}(\mathcal{B})\to C_{c}(\mathcal{L}),$ $f\mapsto\tilde{f}$ with $\tilde{f}(u)=f|_{u}$, extends to a *-isomorphim $\Phi\colon L^{1}(\mathcal{B})\to L^{1}(\mathcal{L})$ (see [17, VIII 6.7]). The cross-sectional $\mathrm{C}^{*}$-algebras of $\mathcal{B}$ and $\mathcal{L}$ are the envelping $\mathrm{C}^{*}$-algebras of $L^{1}(\mathcal{B})$ and $L^{1}(\mathcal{L})$, so we may think $C^{*}(\mathcal{B})=C^{*}(\mathcal{L})$. In addition, the map $\rho\colon\mathcal{L}\to{\mathcal{B}}_{/H}$ of [17, VIII 17.7] is a dense embedding in the sense of [17, XI 12.6], so $\{\rho\circ f\colon f\in C_{c}(\mathcal{L})\}$ is dense in $C_{c}({\mathcal{B}}_{/H})$ with respect to the inductive limit topology (use [17, II 14.6]). The construction of the bundle $\mathrm{C}^{*}$-completion implies that all *-representations of $\mathcal{L}$ have the form $T\circ\rho$ for a *-representation $T$ of ${\mathcal{B}}_{/H}.$ Then the continuous *-homomorphism $L^{1}(\mathcal{L})\to L^{1}({\mathcal{B}}_{/H})$ mapping $f\in C_{c}(\mathcal{L})$ to $\rho\circ f$ induces an isomorphism between $C^{*}(\mathcal{L})$ and $C^{*}({\mathcal{B}}_{/H})$. This yields $C^{*}(\mathcal{B})=C^{*}({\mathcal{B}}_{/H})$ or, more precisely, there exists an isomorphism
@@
We claim that the fiber $\dot{B}_{H}$ of ${\mathcal{B}}_{/H}$ over the unit $H$ of $G/H$ is $C^{*}(\mathcal{B}_{H})$. For the bundle $\mathcal{L}$, [17, XI 11.6] says that $C^{*}(\mathcal{B}_{H})=\dot{B}_{H}$ if and only if every *-representation of $L^{1}(\mathcal{B}_{H})$ is $\mathcal{L}$-positive. Using [17, XI 12.7] we get that this last claim is equivalent to say that every *-representation of $\mathcal{B}_{H}$ is $\mathcal{B}$-positive, which is the case [18, Theorem 2.1].
@@ 5.4. The reduced completion of the partial cross-sectional bundle
We want to construct a reduced version ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ of ${\mathcal{B}}_{/H}$ so that $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is isomophic to $C^{*}_{\mathop{\rm r}}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$ ($H$ is normal and closed in $G$). To do this we take a nondegenerate *-representation $T\colon\mathcal{B}\to\mathbb{B}(X)$ with $\phi:=T|_{B_{e}}$ faithful, and let $S\colon\mathcal{L}\to\mathbb{B}(Y)$ be the *-representation corresponding to $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)$ via the process of [17, VIII 15.9]. Using $S$ and the construction of [17, VIII 16.7] we can produce a Fell bundle ${\mathcal{B}}^{\mathop{\rm r}}_{/H}=\{\dot{B}^{\mathop{\rm r}}_{u}\}_{u\in G/H}$ and a dense embedding $\rho^{\mathop{\rm r}}\colon\mathcal{L}\to{\mathcal{B}}^{\mathop{\rm r}}_{/H}$ such that, for all $f\in\mathcal{L},$ $\|\rho^{\mathop{\rm r}}(f)\|=\|S_{f}\|$. By construction, for all $r\in G$; $f\in C_{c}(\mathcal{B}_{rH})$ and $\xi\in Y,$ $S_{f}\xi=\int_{N}\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)_{f(rt)}\xi dt$. Notice that we may replace $f(rt)$ with $\tilde{f}(rN)(rt)$. This ultimately implies that $\tilde{S}\circ\Phi|_{C_{c}(\mathcal{B})}=\widetilde{\operatorname{Ind}}_{B_{e}\uparrow\mathcal{B}}(\phi)|_{C_{c}(\mathcal{B})}$.
The fiber $\dot{B}^{\mathop{\rm r}}_{H}$ of ${\mathcal{B}}^{\mathop{\rm r}}_{/H}$ over the unit $H$ of $G/H$ is the completion of $L^{1}(\mathcal{B}_{H})$ with respect to the seminorm $f\mapsto\|S_{f}\|.$ Hence, to prove $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ it suffices to show that for all $f\in C_{c}(\mathcal{B}_{H})$ we have $\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})}=\|S_{f}\|.$ By Fell’s absorption principle [16, 18], $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(\phi)\equiv\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}}(T|_{B_{e}})$ is weakly equivalent ($\sim$) to $\lambda^{G}\otimes T.$ The process of restricting *-representations is continuous with respect to the regional topology [17, VIII 21.20] and, by definition, it is a homeomorphism with respect to the operation of taking integrated forms. This implies
@@
where the last equivalence holds because $\lambda^{G}|_{H}\sim\lambda^{H}$ and because the process of taking tensor products with representations of the base group is continuous with respect to the regional topology. By [16, Corollary 2.15], for all $f\in L^{1}(\mathcal{B}_{H})$ we have $\|S_{f}\|=\|\lambda^{H}\tilde{\otimes}(T|_{\mathcal{B}_{H}})(f)\|=\|f\|_{C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})}.$ Then $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$.
@@
The integrated form of $\operatorname{Ind}_{B_{e}\uparrow\mathcal{B}_{H}}(T|_{B_{e}})\sim\lambda^{H}\otimes(T|_{\mathcal{B}_{H}})$ is a *-representation $R$ of $L^{1}(\mathcal{B}_{H})$ for which there exists a faithful *-representation $R^{\prime}$ of $\dot{B}^{\mathop{\rm r}}_{H}=C^{*}_{\mathop{\rm r}}(\mathcal{B}_{H})$ such that $R^{\prime}\circ\rho|_{L^{1}(\mathcal{B}_{H})}=R,$ see [16, Corollary 2.15]. By [17, XI 12.6-12.7], $S=\operatorname{Ind}_{L^{1}(\mathcal{B}_{H})\uparrow\mathcal{L}}(R)$ and $\operatorname{Ind}_{\dot{B}^{\mathop{\rm r}}_{H}\uparrow{\mathcal{B}}^{\mathop{\rm r}}_{/H}}(R^{\prime})\circ\rho^{\mathop{\rm r}}=\operatorname{Ind}_{L^{1}(\mathcal{B}_{H})\uparrow\mathcal{L}}(R).$ From Proposition 3.1 of [18] we get the existence of a faithful *-representation
@@
The function $\Omega^{\mathop{\rm r}}\colon C_{c}(\mathcal{B})\to C_{c}({\mathcal{B}}^{\mathop{\rm r}}_{/H}),$ $f\mapsto\rho^{\mathop{\rm r}}\circ\tilde{f}$, is a *-homomorphism because $\nu\circ\Omega^{\mathop{\rm r}}=\mu$. Using [17, II 14.6] one can show that the image of $\Omega^{\mathop{\rm r}}$ is dense with respect to the inductive limit topology, so $\Omega^{\mathop{\rm r}}(C_{c}(\mathcal{B}))$ is dense in $C^{*}({\mathcal{B}}^{\mathop{\rm r}}_{/H})$ and $\Omega^{\mathop{\rm r}}$ extends to an isomorphism